OpenStudy (anonymous):

Find the common ratio of the sequence: 320, 80, 20, 5, 1.25, …

OpenStudy (anonymous):

1/4

OpenStudy (mathstudent55):

Correct. To find the common ratio, just divide any term (except the first one) by the previous term. For example, 80/320 = 1/4 Also, 5/20 = 1/4